Mang Akong invested a certain amount at 36% annual interest rate, three-halves of this amount at 42%.If his total annual income is 36,900 pesos how much was his total investment?
Answer by macston(5194)

.
T=amount at 36%; F=amount at 42%=1.5T
.
0.36T+0.42F=36900
0.36T+0.42(1.5T)=36900
0.36T+0.63T=36900
0.99T=36900
T=37273
He invested 37273 at 36%.
.
F=1.5T=1.5(37273)=55909
He invested 55909 at 42%.
.
Total Invested=T+F
Total Invested=37273+55909=93182
.
ANSWER: His total investment was 93182 pesos.
.
CHECK:
0.36T+0.42F=36900
0.36(37273)+0.42(55909)=36900
13418+23482=36900
36900=36900
.